# Morphology-Consistent Humanoid Interaction through Robot-Centric Video Synthesis

### 20 Mar 2026 v1

@@ 4.2 Real-World Zero-Shot Interaction
To evaluate morphology-aware interaction, we define four diverse, spatially-sensitive tasks: (1) Kicking a ball; (2) Hugging a box; (3) Punching a bag; and (4) Sitting on a sofa.
@@
Metrics and Results. We report the Success Rate and the Average Spatial Alignment Error (calculated as the Mean Absolute Error of the end-effector distance to the intended contact point) across 20 trials per task. To evaluate robustness over distance, we also report the initial Task Length (the physical distance from the robot’s starting position to the interaction target). As shown in Table 1 and Table 2, the human-centric baseline fundamentally fails on spatial-sensitive tasks. Furthermore, we observe that traditional retargeting errors are severely compounded under locomotion. For instance, in the ’Kick Ball (d)’ task requiring 2.18m of mobility, the human-centric baseline accumulates an overwhelming 0.79m of spatial misalignment. In contrast, Dream2Act effectively bounds the execution error (averaging 0.14m across kicking tasks) regardless of the task length. By operating natively within the G1’s embodiment, Dream2Act preserves precise spatial relationships and absolute metric scales, achieving significantly higher success rates and robust task completion.
@@ 4.3 Text-Driven Free-Space Motion Generation
Figure 6: Qualitative Comparison of Text-to-Motion Pipelines in Free Space. Conventional human-centric T2M models amplify inherent artifacts during retargeting, causing dynamic constraint violations and hardware failure. Dream2Act strictly enforces physical consistency, achieving stable zero-shot deployment.
@@ 4.4 Module Evaluation I: Native Pose Perception
Robust 2D visual perception is the prerequisite for kinematic recovery. We evaluate our 2D detection module on both the Simulation Test Set and the In-the-Wild Internet Test Set.
Baselines and Ablation. We compare against zero-shot human pose estimators (ViTPose [35] pre-trained on COCO). Furthermore, we ablate our training strategy by comparing a variant fine-tuned solely on simulation data (Ours-Sim Only) against our full model trained on the hybrid dataset (Ours-Hybrid).
Results. As detailed in Table 3, off-the-shelf human detectors fail catastrophically on the G1 morphology. While the Sim Only variant performs flawlessly on simulated images, its Average Precision (AP) and 2D PCK drops precipitously on the Internet test set due to lighting and texture domain shifts. Our Hybrid model effectively closes this Sim-to-Real gap, achieving robust native joint localization across all environments.
@@
4.5 Module Evaluation II: Evaluation of 2D-to-3D Lifting
The precision of converting 2D pixel coordinates $\mathbf{k}_{2D}$ into 3D joint positions $\mathbf{P}_{3D,L}$ is fundamental to the fidelity of the final humanoid motion. To quantify the performance of our 2D-to-3D lifting network $\Phi_{L}(\cdot)$, we conduct a rigorous evaluation on the simulation test set, which provides absolute 3D ground truth for the G1 robot $\mathcal{M}_{G1}$. Using Mean Per Joint Position Error (MPJPE) as the primary metric, our lifting network achieves a remarkably low average error of 29mm on the simulation benchmark. This high precision confirms the network’s capability to accurately regress the robot’s native-scale 3D articulation from monocular keypoints, effectively overcoming the depth ambiguity inherent in single-view perception by explicitly modeling G1-specific joint distributions. Such high-fidelity 3D reconstruction provides a geometrically consistent prior for the subsequent URDF-constrained IK solver, ensuring that the recovered trajectories $\tau_{1:T}$ not only visually align with the generative hallucinations but also strictly adhere to the robot’s physical dimensions for stable real-world deployment.
@@ 5 Conclusion
In this paper, we presented Dream2Act, a pioneering robot-centric framework that enables true zero-shot humanoid interaction through generative video synthesis. By shifting from traditional human-centric retargeting to directly hallucinating interactions in the robot’s native spatial coordinate system, our approach fundamentally eliminates the persistent morphology gap and spatial misalignments that have long bottlenecked dexterous skill acquisition. To seamlessly bridge the pixel-to-physics gap, we introduced a robust, domain-specific kinematic recovery pipeline—cascading native 2D pose estimation, 2D-to-3D lifting, and URDF-constrained inverse kinematics—to translate visual priors into physically viable joint trajectories.Extensive physical validations on the Unitree G1 humanoid across a diverse set of spatially-sensitive interactions (e.g., ball kicking, sofa sitting, and bag punching) demonstrate that Dream2Act reliably achieves correct physical contacts and significantly outperforms conventional optimization-based baselines in both spatial alignment and task success rates. While currently functioning as an offline zero-shot motion planner due to the inherent inference latency of diffusion models , Dream2Act bypasses prohibitive data collection and task-specific policy training, establishing a highly scalable, data-driven pathway for advancing embodied humanoid intelligence.

### 24 Mar 2026 v2

@@ 4.2 Real-World Zero-Shot Interaction
To evaluate morphology-aware interaction, we define four diverse, spatially-sensitive tasks: (1) Ball-Kicking; (2) Box-hugging; (3) Bag-hitting; and (4) Sofa-sitting.
@@
Metrics and Results. We report the Success Rate and the Average Spatial Alignment Error (calculated as the Mean Absolute Error of the end-effector distance to the intended contact point) across 20 trials per task. To evaluate robustness over distance, we also report the initial Task Length (the physical distance from the robot’s starting position to the interaction target). As shown in Table 1 and Table 2, the human-centric baseline fundamentally fails on spatially-sensitive tasks. Furthermore, we observe that traditional retargeting errors are severely compounded under locomotion. For instance, in the ’Kick Ball (d)’ task requiring 2.18m of mobility, the human-centric baseline accumulates an overwhelming 0.79m of spatial misalignment. In contrast, Dream2Act effectively bounds the execution error (averaging 0.14m across kicking tasks) regardless of the task length. By operating natively within the G1’s embodiment, Dream2Act preserves precise spatial relationships and absolute metric scales, achieving significantly higher success rates and robust task completion.
@@ 4.3 Text-Driven Free-Space Motion Generation
Figure 6: Qualitative Comparison of Text-to-Motion Pipelines in Free Space. Human-centric T2M models amplify inherent artifacts during retargeting, causing dynamic constraint violations and hardware failure. Dream2Act strictly enforces physical consistency, achieving stable zero-shot deployment.
@@ 4.4 Module Evaluation I: Native Pose Perception
Robust 2D perception is the prerequisite for kinematic recovery. We evaluate our 2D detection module on both the Simulation Test Set and the In-the-Wild Internet Test Set. Baselines and Ablation. We compare against zero-shot human pose estimators (ViTPose [35] pre-trained on COCO). We ablate our training strategy by comparing a variant fine-tuned solely on simulation data (Ours-Sim Only) against our model trained on the hybrid dataset (Ours-Hybrid).
Results. As detailed in Table 3, off-the-shelf human detectors fail catastrophically on the G1 morphology. While the Sim Only variant performs flawlessly on simulated images, its Average Precision (AP) and 2D PCK drop precipitously on the Internet test set due to lighting and texture domain shifts. Our Hybrid model effectively closes this Sim-to-Real gap, achieving robust native joint localization across all environments.
@@
4.5 Module Evaluation II: 2D-to-3D Lifting
The precision of converting 2D pixel coordinates $\mathbf{k}_{2D}$ into 3D joint positions $\mathbf{P}_{3D,L}$ is fundamental to the fidelity of the final humanoid motion. To quantify the performance of our 2D-to-3D lifting network $\Phi_{L}(\cdot)$, we conduct a rigorous evaluation on the simulation test set, which provides absolute 3D ground truth for the G1 robot $\mathcal{M}_{G1}$. Using Mean Per Joint Position Error (MPJPE) as the primary metric, our lifting network achieves a low average error of 29mm on the simulation benchmark. This high precision confirms the network’s capability to accurately regress the robot’s native-scale 3D articulation from monocular keypoints, effectively overcoming the depth ambiguity inherent in single-view perception by explicitly modeling G1-specific joint distributions. Such high-fidelity 3D reconstruction provides a geometrically consistent prior for the subsequent URDF-constrained IK solver, ensuring that the recovered trajectories $\tau_{1:T}$ not only visually align with the generative hallucinations but also strictly adhere to the robot’s physical dimensions for stable real-world deployment.
@@ 5 Conclusion
In this paper, we presented Dream2Act, a pioneering robot-centric framework that enables true zero-shot humanoid interaction through generative video synthesis. By shifting from traditional human-centric retargeting to directly synthesizing interactions in the robot’s native spatial coordinate system, our approach fundamentally eliminates the persistent morphological gap and spatial misalignments that have long bottlenecked dexterous skill acquisition. To seamlessly bridge the pixel-to-physics gap, we introduced a robust, domain-specific kinematic recovery pipeline—cascading native 2D pose estimation, 2D-to-3D lifting, and URDF-constrained inverse kinematics—to translate visual priors into physically viable joint trajectories. Extensive physical validations on the Unitree G1 humanoid across a diverse set of spatially-sensitive interactions (e.g., ball kicking, sofa sitting, and bag punching) demonstrate that Dream2Act reliably achieves correct physical contacts and significantly outperforms conventional optimization-based baselines in both spatial alignment and task success rates. While currently functioning as an offline zero-shot motion planner due to the inherent inference latency of diffusion models, Dream2Act bypasses prohibitive data collection and task-specific policy training, establishing a highly scalable, data-driven pathway for advancing embodied humanoid intelligence, with future work focusing on accelerating inference for real-time closed-loop control.